The area of a rectangle is 72 centimeters^2. What is the value of n if the length is n+9 and the width is .4?
I've tried and gotten two complete different answers and when checked they aren't even close to the answer. I've tried 72=n+9/.4 and got 19.18 which when checked using the original formula i got 16.92 and that is nothing close to 72. I tried doing it like this 72=.4(n+9) and got 62.6 and when checked i got 34.04 which is also not correct. I need help soon. I've asked friends also and their methods don't work.

Answer by ankor@dixie-net.com(22740)

The area of a rectangle is 72 centimeters^2.
What is the value of n if the length is n+9 and the width is .4?
:
Area is length * width, therefore:
.4(n+9) = 72
.4n + 3.6 = 72
.4n = 72 - 3.6
.4n = 68.4
n = 68.4/.4
n = 171 cm
:
:
See if that checks out
L = 171 + 9 = 180
180 * .4 = 72